# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 3. Reduced clique-cube complexes and actions on them
$\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})}$
$w_{\mathfrak{j}(w)}(h-1)\cdots w_{\mathfrak{j}(w)}(1)w_{\mathfrak{j}(w)}(0)A_{V_{0}}$
$=\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})}$
$w(h-1)\cdots w(1)w(0)A_{V_{0}}$
where we note $w(h-1)=w_{\mathfrak{j}(w)}$ and $w(0)=v_{i(w),l(w)}$. Hence
(4.11) $\begin{split}g\in&\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w))\\ &A_{V_{0}}(\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w)))^{-1}.\end{split}$
Since $g$ commutes with
$\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}$
$v_{1,l(w_{\mathfrak{j}(w)})}\cdots v_{i(w_{\mathfrak{j}(w)})-1,l(w_{\mathfrak{j}(w)})}v_{i(w_{\mathfrak{j}(w)})+1,l(w_{\mathfrak{j}(w)})}\cdots v_{k,l(w_{\mathfrak{j}(w)})},$
$g\in(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$. By noting $w(0)<w(1)<\cdots<w(h-1)<w(h)=w$, we can show $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}>w\}}$. Indeed it follows from Lemma 2.11 that $g$ belongs to
$A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap A_{V_{0}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap(w(h-1)\cdots w(1))A_{V_{0}}(w(h-1)\cdots w(1))^{-1}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap A_{V_{0}\setminus\{w(1)\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2))^{-1}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}>w\}}.$
Since $<$ is the total order on $V_{0}=W(1)\sqcup\cdots\sqcup W(\mathfrak{d})$, which is a finite set, we repeat the same argument and have $g=1$. ∎
@@
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then, Theorem 1.2 $(1)\Rightarrow(2)$ implies that $A_{\Gamma}$ is acylindrically hyperbolic. Therefore, the center of $A_{\Gamma}$ is finite by acylindrical hyperbolicity ([33, Corollary 7.3]). Proposition 3.8 then implies Corollary 1.4. We can show Corollary 1.5 by a similar way. ∎
@@
There are some versions of the so-called center conjecture. Note that the center of any irreducible Artin group of spherical type is known to be infinite cyclic [4], [17]. Then as a version, the center conjecture containing this known fact can be stated as follows: the center of any irreducible Artin group of spherical type and of infinite type is infinite cyclic and trivial, respectively. Since the center of any possibly reducible Artin group with each irreducible factor satisfying the center conjecture is trivial or torsion-free, the last part of Corollary 1.4 obviously implies the following: in the setting of Corollary 1.4, if each irreducible factor of $A_{\Gamma_{0}}$ satisfies the center conjecture, then $A_{\Gamma}$ satisfies the center conjecture. This claim is independently given in [27] by a different way.

### v3

@@ 3. Reduced clique-cube complexes and actions on them
$\lambda_{l_{\sharp}(w)}=v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)}\cdots v_{k,l_{\sharp}(w)}.$
Since the action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$ preserves the cube structure, all cubes penetrated by $\ell$ are fixed by $g$. Therefore, $g$ fixes
$\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}},$
see Figure 10. Then
$g\in(\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)})A_{V_{0}}(\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w),l_{\sharp}(w)})^{-1}.$
By $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$ and Lemma 4.3 (1), $\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}v_{1,l_{\sharp}(w)}\cdots v_{i_{\sharp}(w)-1,l_{\sharp}(w)}$ commutes with $g$, see Figure 7. Therefore, $g\in v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}$, and thus $g\in A_{V_{0}}\cap v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}$. According to Lemma 2.11, $A_{V_{0}}\cap v_{i_{\sharp}(w),l_{\sharp}(w)}A_{V_{0}}v_{i_{\sharp}(w),l_{\sharp}(w)}^{-1}\subset A_{V_{0}\setminus\{w\}}$. Hence
$g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}\setminus\{w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}.$
(iii-3-$b$) We consider the case where $w\in W(h)$ where $h\geq 2$. Then we have
$(w(h),w(h-1),\ldots w(0))=(w,\pi(w),\ldots\pi^{h}(w))$
[…]
(4.36) $\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\lambda_{\sharp}(\mathfrak{j})A_{V_{0}},$
which is fixed by $g$. We note that $\lambda_{\sharp}(\mathfrak{j})$ equals
$v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w_{\mathfrak{j}}(h-1)\cdots w_{\mathfrak{j}}(1)w_{\mathfrak{j}}(0)$
$=v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w(h-1)\cdots w(1)w(0).$
Hence
$\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\lambda_{\sharp}(\mathfrak{j})$
$=$ $\left(\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}\right)\cdot$
$\left(v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}w(h-1)\cdots w(1)w(0)\right).$
Note that $\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)$ is a product of elements of $V_{\ast}\sqcup\{w_{1},\ldots w_{\mathfrak{j}-1}\}$. By Lemma 4.3 (2) with $w_{1}\prec\cdots\prec w_{\mathfrak{j}-1}\prec w_{\mathfrak{j}}=\pi(w)$, any $w^{\prime\prime}\succeq w$ commutes with
$\gamma_{\natural}\gamma_{\sharp}(1)\cdots\gamma_{\sharp}(\mathfrak{j}-1)\lambda_{1}\cdots\lambda_{l_{\sharp}(w_{\mathfrak{j}})-1}v_{1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{i_{\sharp}(w_{\mathfrak{j}})-1,l_{\sharp}(w_{\mathfrak{j}})}v_{i_{\sharp}(w_{\mathfrak{j}})+1,l_{\sharp}(w_{\mathfrak{j}})}\cdots v_{k,l_{\sharp}(w_{\mathfrak{j}})}.$
Hence so does $g$. Then $g\in(w(h-1)\cdots w(0))A_{V_{0}}(w(h-1)\cdots w(0))^{-1}$. Next, by noting $w(0)\prec w(1)\prec\cdots\prec w(h-1)\prec w(h)=w$, we can show $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}$. Indeed, it follows from Lemma 2.11 that $g$ belongs to
$A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap(w(h-1)\cdots w(1))A_{V_{0}}(w(h-1)\cdots w(1))^{-1}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap A_{V_{0}\setminus\{w(1)\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}\cap(w(h-1)\cdots w(2))A_{V_{0}\setminus\{w(1)\}}(w(h-1)\cdots w(2))^{-1}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$\subset$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$
@@
$=$ $A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}.$
In the above deformation, all $\subset$ are consequences of Lemma 2.11, and the final equality follows from $w(h)=w$. The other equalities are trivial deformations.
Since $\prec$ is a total order on a finite set $V_{0}$, by applying (iii-3-$a$) or (iii-3-$b$) for $w\in V_{0}$ in order from the smallest element of $V_{0}$ to the largest one, we have $g\in A_{\emptyset}=\{e\}$ and thus $g=e$. ∎
@@
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then, Theorem 1.2 implies that $A_{\Gamma}$ is acylindrically hyperbolic. Therefore, the center of $A_{\Gamma}$ is finite by acylindrical hyperbolicity ([33, Corollary 7.3]). Proposition 3.8 then implies Corollary 1.4. We can show Corollary 1.5 in a similar way. ∎
@@
There are several versions of the so-called center conjecture. Note that the center of any irreducible Artin group of spherical type is known to be infinite cyclic [4], [17]. Then, as one version, the center conjecture containing this known fact can be stated as follows: the center of any irreducible Artin group of spherical type and of infinite type is infinite cyclic and trivial, respectively. Since the center of any possibly reducible Artin group with each irreducible factor satisfying the center conjecture is trivial or torsion-free, the last part of Corollary 1.4 clearly implies the following: in the setting of Corollary 1.4, if each irreducible factor of $A_{\Gamma_{0}}$ satisfies the center conjecture, then $A_{\Gamma}$ satisfies the center conjecture. This claim is independently given in [27] by a different method.
5. Example
Let $\Gamma$ be the graph with $V(\Gamma)=\{x,y,u_{1},u_{2},v_{1},v_{2}\}$ as shown in Figure 1. In this section, we explain the construction of $\gamma_{\natural\sharp\flat}$ for this example.
Remark 5.1.
Theorems 1.2 and 4.1 can be applied to Artin groups associated with $\Gamma$ in Figures 4 and 5. However, these examples may be too simple to illustrate the construction of $\gamma_{\natural\sharp\flat}$. Indeed, in the leftmost defining graph in Figure 4, $\gamma_{\flat}=e\in A_{\Gamma}$. For the middle and right $\Gamma$, $\gamma_{\sharp}=\gamma_{\flat}=e\in A_{\Gamma}$. Also, for all $\Gamma$ in Figure 5, $\gamma_{\sharp}=e\in A_{\Gamma}$. The graph $\Gamma$ in Figure 1 is one of the smallest that has non-trivial $\gamma_{\sharp}$ and $\gamma_{\flat}$.
We note that $\Gamma$ is irreducible; that is, $\Gamma^{t}$ is connected. Then we have the join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast(\Gamma_{1}\ast\Gamma_{2})$, where we set subgraphs $\Gamma_{i}$ $(i=0,1,2)$ of $\Gamma$ by
(5.1) $V_{0}:=V(\Gamma_{0})=\{x,y\},\ E_{0}:=E(\Gamma_{0})=\{(x,y),\ (y,x)\},$
(5.2) $V_{1}:=V(\Gamma_{1})=\{u_{1},u_{2}\},\ E_{1}:=E(\Gamma_{1})=\emptyset,$
(5.3) $V_{2}:=V(\Gamma_{2})=\{v_{1},v_{2}\},\ E_{2}:=E(\Gamma_{2})=\emptyset.$
Clearly, $\Gamma_{0}$ is the maximal clique factor, and $\Gamma_{1}$ and $\Gamma_{2}$ are join-indecomposable factors. Note that $k=2$. We give an explicit construction of $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}$ for this $\Gamma$ following Section 4.
First, we construct $\gamma_{\natural}$. In our $\Gamma$, $Q(\Gamma,\Gamma_{0})$ is the tree consisting of one edge and two vertices $V_{1}$ and $V_{2}$ (see (LABEL:QGamma)). Here, the spanning tree $T$ coincides with $Q(\Gamma,\Gamma_{0})$. We choose $V_{1}$ as the root of $T$. Take paths $p_{1,2}$ and $p_{2,1}$ in $\Gamma^{t}$ as follows:
(5.4) $p_{1,2}=(w_{1,2,0},w_{1,2,1},w_{1,2,2})=(u_{1},x,v_{1}),$
(5.5) $p_{2,1}=(w_{2,1,0},w_{2,1,1},w_{2,1,2})=(v_{1},x,u_{1})$
(see (4.2) and (4.3)). Their end vertices are
(5.6) $s_{1,2}=w_{1,2,0}=u_{1}\in V_{1},\ t_{1,2}=w_{1,2,2}=v_{1}\in V_{2},$
(5.7) $s_{2,1}=w_{2,1,0}=v_{1}\in V_{2},\ t_{2,1}=w_{2,1,2}=u_{1}\in V_{1},$
respectively (see (4.4)). Set $\tau_{1,2},\tau_{2,1}\in A_{\Gamma}$ as
(5.8) $\tau_{1,2}=w_{1,2,0}w_{1,2,1}w_{1,2,2}=u_{1}xv_{1},$
(5.9) $\tau_{2,1}=w_{2,1,0}w_{2,1,1}w_{2,1,2}=v_{1}xu_{1}$
(see (4.5)). We fix a closed path on $T=Q(\Gamma,\Gamma_{0})$;
(5.10) $(V_{i_{1}},V_{i_{2}},V_{i_{3}})=(V_{1},V_{2},V_{1})$
(see (4.6)). We note that the length $r$ of the path is $2$.
According to Lemma 4.2, we construct $n\in\mathbb{N}$, $M_{\natural}\in\mathcal{M}(2,n;V(\Gamma))$ and a map $l_{\flat}$. Take closed paths of length $n=2$ on $\Gamma_{1}^{c}$, $\Gamma_{2}^{c}$:
(5.11) $(u_{1},u_{2},u_{1}),\ (v_{1},v_{2},v_{1}),respectively.$
Set
(5.12) $M_{\natural}=\begin{pmatrix}u_{1}&u_{2}\\ v_{1}&v_{2}\end{pmatrix},$
and
(5.13) $l_{\flat}(1,2)=l_{\flat}(2,1)=1.$
Then,
(5.14) $M_{\natural}[1,l_{\flat}(1,2)]=u_{1},\ M_{\natural}[2,l_{\flat}(2,1)]=v_{1}.$
We have
(5.15) $\gamma_{\natural}=u_{1}v_{1}u_{2}v_{2},$
[…]
(5.16) $M^{\prime}_{\flat}(1)=M^{\prime}_{\flat}(2)=\begin{pmatrix}u_{1}^{-1}&e\\ v_{1}^{-1}&e\end{pmatrix},$
[…]
(5.17) $M^{\prime\prime}_{\flat}(1)=\begin{pmatrix}\tau_{1,2}&e\\ e&e\end{pmatrix}=\begin{pmatrix}u_{1}xv_{1}&e\\ e&e\end{pmatrix}$
and
(5.18) $M^{\prime\prime}_{\flat}(2)=\begin{pmatrix}\tau_{2,1}&e\\ e&e\end{pmatrix}=\begin{pmatrix}v_{1}xu_{1}&e\\ e&e\end{pmatrix},$
[…]
(5.19) $M_{\flat}(1)=\begin{pmatrix}u_{1}xv_{1}&u_{2}\\ e&v_{2}\end{pmatrix},$
and
(5.20) $M_{\flat}(2)=\begin{pmatrix}v_{1}xu_{1}&u_{2}\\ e&v_{2}\end{pmatrix},$
[…]
(5.21) $M_{\flat}=\begin{pmatrix}u_{1}xv_{1}&u_{2}&v_{1}xu_{1}&u_{2}\\ e&v_{2}&e&v_{2}\end{pmatrix},$
[…]
(5.22) $\gamma_{\flat}=(u_{1}xv_{1})(u_{2}v_{2})(v_{1}xu_{1})(u_{2}v_{2}),$
[…]
(5.23) $W(1)=\{x\},\ W(2)=\{y\}.$
Note that $d_{\Gamma^{t}}(x,V_{\ast})=1$, $d_{\Gamma^{t}}(y,V_{\ast})=2$. Then $V_{0}=W(1)\sqcup W(2)$ and $\mathfrak{h}=2$. Here, $d_{\Gamma^{t}}(x,V_{\ast})=d_{\Gamma^{t}}(x,u_{1})=d_{\Gamma^{t}}(x,v_{1})$. Therefore,
(5.24) $(i_{\sharp}(x),l_{\sharp}(x))=(i_{\sharp}(y),l_{\sharp}(y))=(1,1)$
[…]
(5.25) $w_{1}=w_{\mathfrak{m}}=x$
where
(5.26) $\mathfrak{m}=\#\left(W(1)\sqcup\cdots\sqcup W(\mathfrak{h}-1)\right)=\#W(1)=1$
[…]
(5.27) $l_{\sharp}(w_{1})=l_{\sharp}(x)=1.$
[…]
(5.28) $M^{\prime}_{\sharp}(1)=\begin{pmatrix}u_{1}^{-1}&e\\ e&e\end{pmatrix},$
[…]
(5.29) $M^{\prime\prime}_{\sharp}(1)=\begin{pmatrix}e&e\\ w_{1}(1)w_{1}(0)&e\end{pmatrix}=\begin{pmatrix}e&e\\ xu_{1}&e\end{pmatrix},$
[…]
(5.30) $M_{\sharp}(1)=\begin{pmatrix}e&u_{2}\\ v_{1}xu_{1}&v_{2}\end{pmatrix},$
[…]
(5.31) $M_{\sharp}=M_{\sharp}(1)=\begin{pmatrix}e&u_{2}\\ v_{1}xu_{1}&v_{2}\end{pmatrix},$
[…]
(5.32) $\gamma_{\sharp}=(v_{1}xu_{1})(u_{2}v_{2}),$
[…]
Then
(5.33) $\gamma_{\natural\sharp\flat}=\gamma_{\natural}\gamma_{\sharp}\gamma_{\flat}=(u_{1}v_{1}u_{2}v_{2})(v_{1}xu_{1}u_{2}v_{2})(u_{1}xv_{1}u_{2}v_{2}v_{1}xu_{1}u_{2}v_{2})$
[…]